# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v3

@@ Appendix A HyperParameters and Training Cost
Figure 5: An overview of our dataset diversification step. It is recommended to view the details with colors.
Appendix B Dataset Diversification
@@ Appendix D Entire fine-tuning results
Appendix E Specific ablation discussion on BBH
@@
Appendix F Detailed results of 6 BBH subsets
  • Overall Performance. The FIPO optimizer performs well in most combinations, usually leading in each task and average score. PromptAgent and GPT-4 optimizers perform well on some specific tasks, but are slightly inferior to FIPO in overall average scores.
  • Task-specific Performance. On a task level, FIPO excels in many tasks. For instance, in the "Temporal" and "Causal Judgment" tasks under Llama2-7B, it achieved the highest scores of 0.683 and 0.616, respectively. It also secured the top scores of 0.525 and 0.891 in the "Penguins" and "Causal Judgment" tasks under Tulu2-7B. For GPT-4o, FIPO achieved the highest scores in the "Geometry" and "Causal Judgment" tasks, with 0.810 and 0.840, respectively. Similarly, in the "Epistemic" task under Qwen2-72B-Instruct, FIPO led with a score of 0.792. Overall, FIPO optimizer demonstrates balanced and strong performance across various tasks.

### v4

@@ Appendix C Additional details of IPL
Appendix D Entire fine-tuning results
Appendix E Detailed results of 6 BBH subsets
Overall Performance. The FIPO optimizer performs well in most combinations, usually leading in each task and average score. PromptAgent and GPT-4 optimizers perform well on some specific tasks, but are slightly inferior to FIPO overall.
Task-specific Performance. FIPO excels in many tasks. In the "Temporal" and "Causal Judgment" tasks under Llama2-7B, it achieved the highest scores of 0.683 and 0.616, respectively. It also secured the top scores of 0.525 and 0.891 in the "Penguins" and "Causal Judgment" tasks under Tulu2-7B. For GPT-4o, FIPO achieved the highest scores in the "Geometry" and "Causal Judgment" tasks, with 0.810 and 0.840, respectively. Similarly, in the "Epistemic" task under Qwen2-72B-Instruct, FIPO led with a score of 0.792.
Appendix F Specific ablation discussion on BBH